OpenStudy (anonymous):

How do you solve f(g(x)) if f(x)=2x^2+x and g(x)=x-2 ? I tried and got 4x^2-15x+18 not sure if I'm right?

OpenStudy (shamim):

\[f(g(x))=f(x-2)=2(x-2)^{2}+x-2=2(x ^{2}-4x+4)+x-2=2x ^{2}-8x+8+x-2=2x ^{2-7x+6}\]

OpenStudy (shamim):

did u catch my solution

OpenStudy (anonymous):

yes, but when you got 2(x^2−4x+4)+x−2 how'd you get your final answer?

OpenStudy (anonymous):

for my final answer I got 2x^2-7x+6 is that correct now?

OpenStudy (anonymous):

if f(x)=2x^2+x and g(x)=x-2 ----------------- plug in (x-2) into every x in 2x^2+x so 2(x-2)^2+(x-2) ^ might wanna distribute the 2 here. and and get (2x-4)^2 but you not don yet cuzz (2x-4)^2 can also be expanded (2x-4) * (2x-4) after that combine like terms :D

OpenStudy (anonymous):

here it is expanded fully \[2\,{x}^{2}-7\,x+6\]

OpenStudy (anonymous):

Okay so I got it right! Thanks you guys!

OpenStudy (anonymous):

\[2\,{x}^{2}-7\,x+6\] Is the right answer i used maple to dbl check.

OpenStudy (anonymous):

Yeah :) I wrote it above you message!

OpenStudy (anonymous):

YEss good job :D

OpenStudy (anonymous):

Thank You good job too :)!

OpenStudy (anonymous):

ooh btw do you do your math on paper or the pc? if you do it by pc the easyest way is to copy and paste. And either way you do it use lots of perenthesis :D dont be afraid to abuse them.